OpenStudy (goformit100):

Find the smallest natural number which is 4 times smaller than the number written with the same digit but in the reverse order.

OpenStudy (goformit100):

@Thoughts

OpenStudy (anonymous):

Give me some time, please!

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

Confused. @Zarkon @ganeshie8 @amistre64

OpenStudy (anonymous):

[Sorry for tagging you mods but I think you all may help here. ]

OpenStudy (amistre64):

reverse order would suggest at least a 2 digit number

OpenStudy (amistre64):

if we let x equal 10 .... (bx + a)/4 = ax + b

OpenStudy (amistre64):

if there is not value for this, try it as an x^2 format

OpenStudy (amistre64):

well ... (10b + a)/4 = 10a + b :) 10b + a = 40a + 4b 6b = 41a a/b = 6/41

OpenStudy (amistre64):

6/39 that is ... forgot how to subtract "1"

OpenStudy (amistre64):

reducing gets us 2/13 which suggest to me that this is perhaps a 3 digit number at best

OpenStudy (anonymous):

The new number = 5 (the original one)

OpenStudy (amistre64):

100c + 10b + a = 400a +40b + 4c

OpenStudy (anonymous):

If we take :The new number = 4 * the original one Then we have 3 times smaller not 4 times.

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

So, if I am not wrong it should be : The new number = 5 * the original number

OpenStudy (anonymous):

Do you agree amistre?

OpenStudy (amistre64):

4 times bigger than x is: 4x 4 times smaller than x is: x/4 not too sure how you are seeing this tho

OpenStudy (goformit100):

ok ... ... ...

OpenStudy (amistre64):

if we go with my original a:b setup; we have a 2:13 ratio such that 10(13) + 2 = 40(2) + 4(13) 132/4 = 10(2) + 13 132/4 = 33 which doesnt really fit since: a or b > 9 there isnt a nice reversal

OpenStudy (goformit100):

Ok thanks @amistre64

OpenStudy (amistre64):

multiples of 4 end in: 0,4,8,2,6 4n > 9 when n=3 4n > 99 when n = 25 4n > 999 when n = 250 4n > 9999 when n = 2500 at best, we want a number that starts and ends with a 2

OpenStudy (amistre64):

not quite what i mean; the ends of the number are 2 and 4 2xxx4 and 4xxx2

OpenStudy (amistre64):

hmm, make that 2 and 8; since 2/8 is about 1/4 and not 2/4

OpenStudy (anonymous):

I

OpenStudy (anonymous):

I tried a lot , amistre but there was no such number.

OpenStudy (anonymous):

Awesome @amistre64 , it is 2718 * 4 = 8712

OpenStudy (anonymous):

It was sure that the last and the unit digits must be 2 or 8. So, 2..8 * 4 = 8..2 Using modulo further I got. (The hundreds digit) mod 10 = 4(The tens digit) + 3 hmm, that means the hundreds digit is 1 and the tens digit has to be 2 or 7 So, the no. can be : 2128 or 2178 2128 * 4 = 8512 (So this is not true) 2178 * 4 = 8712 (its true now)

OpenStudy (amistre64):

good job :)

OpenStudy (anonymous):

Thanks a lot amistre.